# ChartMoE: Mixture of Expert Connector for Advanced Chart Understanding

### v1

@@ Abstract
Automatic chart understanding is crucial for content comprehension and document parsing. Multimodal large language models (MLLMs) have demonstrated remarkable capabilities in chart understanding through domain-specific alignment and fine-tuning. However, the application of alignment training within the chart domain is still underexplored. To address it, we propose ChartMoE, which employs the mixture of expert (MoE) architecture to replace the traditional linear projector to bridge the modality gap. Specifically, we train multiple linear connectors through distinct alignment tasks, which are utilized as the foundational initialization parameters for different experts. Additionally, we introduce ChartMoE-Align, a dataset with over 900K chart-table-JSON-code quadruples to conduct three alignment tasks (chart-table/JSON/code). Combined with the vanilla connector, we initialize different experts in four distinct ways and adopt high-quality knowledge learning to further refine the MoE connector and LLM parameters. Extensive experiments demonstrate the effectiveness of the MoE connector and our initialization strategy, e.g., ChartMoE improves the accuracy of the previous state-of-the-art from 80.48% to 84.64% on the ChartQA benchmark.
@@ 1 Introduction
Charts serve as a fundamental tool for data visualization, with automated chart interpretation gaining prominence in domains such as text analysis Hoque et al. (2017), scientific research Hsu et al. (2021), and policy-making Wu et al. (2024). Chart understanding is a complex task that demands the identification of visual cues, the comprehension of intricate interactions, and the precise inference of values informed by prior knowledge. Previous work Lee et al. (2023); Liu et al. (2023b, a) typically pretrained on structured datasets or domain-specific charts to improve the model performance and efficiency. However, they are constrained by limited resources and narrow task focus, thereby restricting the application in diverse real-world scenarios. In contrast, MLLMs Li et al. (2023); Liu et al. (2023d); Bai et al. (2023a); Ye et al. (2023b); Chen et al. (2023a); OpenAI (2023) exhibit substantial potential in image comprehension and instruction following. With the exponential growth of chart data, automated chart interpretation via MLLMs is emerging as a promising avenue.
The community has achieved advanced progress by creating chart understanding datasets Liu et al. (2023c); Han et al. (2023); Masry et al. (2024b); Xu et al. (2023) and applying supervised fine-tuning on charts Meng et al. (2024); Yan et al. (2024); Carbune et al. (2024). Recent studies advocate for alignment pre-training as a foundational step in this paradigm, which only updates the connector’s parameters to bridge the modality gap. ChartAst Meng et al. (2024) proposes to align chart with table, significantly improving the performance across several chart benchmarks. ChartReformer Yan et al. (2024) proposes to switch the alignment intermediary from table to information-rich JSON. Some chart MLLMs, e.g., ChartLlama Han et al. (2023), employ code as an intermediary to achieve more accurate chart parsing and precise editing.
@@
Figure 1: Overview and capabilities of ChartMoE: We introduce a MoE architecture connector and provide visualizations of the top-1 expert selection (refer to Fig. 6 and appendix for details). ChartMoE can extract highly precise values and provide flexible chart editing through code-based interactions.
However, few studies have explored how to select appropriate alignment tasks for chart understanding models. For instance, DocOwl Hu et al. (2024) advocates aligning all modalities with markdown, while Llama-v3.1 Team et al. (2024b) warns that excessive markdown data in alignment tasks could be detrimental to model performance. In the chart domain, table and JSON formats primarily capture attributes like titles, axes, and numerical values but are less effective in conveying visual elements such as color and topology. This limitation is especially evident in charts with implicit numerical information, such as network diagrams, where comprehension depends on the spatial relationships between visual elements. In such cases, code effectively captures the relationships between visual marks. Hence, we aim to embrace all these structured text formats.
We introduce ChartMoE, an instruction-tuned MLLM specifically designed for chart comprehension and reasoning. As illustrated in Fig. 1&2, we replace the linear connector with the Mixture of Experts Zoph et al. (2022) (MoE) architecture, each expert utilizes a two-layer MLP with GeLU activation Lee (2023). Specifically, We train three distinct connectors with different alignment tasks independently, i.e., chart to table, chart to JSON, and chart to code, as shown in Fig. 1. For better alignment, we construct ChartMoE-Align via existing chart datasets, which consist of over 900K quadruplets {chart, table, JSON, code}, as illustrated in Fig. 3. Additionally, we retain the original vanilla connector to preserve the base model’s foundational knowledge and conversational abilities. Consequently, we offer four distinct pathways for initializing the expert parameters. Compared to random initialization, ChartMoE achieves faster and more stable convergence. Unlike Co-Upcycling initialization Komatsuzaki et al. (2023), ChartMoE is grounded in a more principled design to prevent experts’ homogenization effectively.
We conduct high-quality knowledge learning using the MMC instruction Liu et al. (2023c) to update the routing network, expert connectors, and LLM LoRA parameters while further employing annealing training with datasets like ChartQA Masry et al. (2022) and ChartGemma Masry et al. (2024b). Following the previous work Zhang et al. (2024); Masry et al. (2024b), we integrate the Program of Thought (PoT) prompting Chen et al. (2023b) to enhance the mathematical computation capabilities. ChartMoE can be deployed for inference on a single A100-40G GPU, achieving state-of-the-art results in areas such as question-answering, number extraction, and chart editing&highlighting. In summary, our contributions are:
@@
    We present ChartMoE for faithful and reasonable chart understanding, with an MoE architecture connector to bridge the visual and LLM branches, enabling differential expert initialization based on various alignment training tasks.
@@
    We introduce ChartMoE-Align, a large-scale dataset with over 900K chart-table-JSON-code quadruplets for chart alignment pretraining.
@@
    We propose to train ChartMoE with a three-stage training paradigm, including alignment pre-training, high-quality knowledge learning, and annealing chart tuning.
@@ 2 Related Work
Multimodal large language models leverages a connector to bridge the gap between large language models Touvron et al. (2023); Radford et al. (2018); Brown et al. (2020); Zhang et al. (2022); Zheng et al. (2023) and vision encoders Radford et al. (2021); Oquab et al. (2023) to enable enriched capabilities of comprehension and instruction following. Approaches such as BLIP2 Li et al. (2023), Flamingo Alayrac et al. (2022), mPLUG-Owl Ye et al. (2023b), and Qwen-VL Bai et al. (2023b) utilize QFormers or Resamplers to align modalities on extensive datasets of image-text pairs. LLaVA Liu et al. (2023d, 2024b) is the pioneering work to extend the instruction tuning paradigm to visual tasks with text-only GPT-4 OpenAI (2023), achieving tremendous performance using a simple MLP without compromising visual information to refine the multimodal alignment. Some works Lin et al. (2023); Tong et al. (2024b, a) explore the combination of various vision encoders, complementarily enhancing visual representations to bolster the fine-grained visual perception of MLLMs. Despite efforts in structural design, training strategies and data quality remain crucial in the advancement of MLLMs.
Chart Reasoning refers to chart analysis, summarization, and etc. Existing methods can be categorized as 1) Two-stage methods use specialized extraction modules to generate intermediary representations of chart information, like tables, which are provided as textual prompts for LLMs. Pix2Struct Lee et al. (2023) aligns markdown data with charts. MatCha Liu et al. (2023b) aligns various data formats (e.g., tables and code) with charts on several downstream tasks. DePlot Liu et al. (2023a) fine-tunes Pix2Struct for table extraction and uses LLMs to process queries based on the extracted data. ChartVLM Xia et al. (2024) employs a discriminator to ascertain the necessity of intervention by LLMs for a given query. 2) End-to-end methods strive to tackle chart reasoning challenges with a unified model. ChartLlama Han et al. (2023) incorporates diverse charts and downstream tasks based on LLaVA Liu et al. (2023d). ChartPaLI Carbune et al. (2024), ChartAst Meng et al. (2024), and MMC Liu et al. (2023c) conduct alignment on table-chart pairs. UReader Ye et al. (2023a) aligns all data with markdown, while mPLUG-Owl2 Ye et al. (2023c) achieves superior performance with high-resolution inputs. ChartThinker Liu et al. (2024c) and DOMINO Wang et al. (2023) propose the CoT Wei et al. (2022) for chart reasoning. LaMenDa Zhuowan et al. (2024) trains MLLMs via step-by-step reasoning QA. ChartReformer Yan et al. (2024) introduces chart-JSON alignment, while OneChart Chen et al. (2024) aligns charts with Python dictionaries. MiniGPT-v2 Chen et al. (2023a), Doc-Owl Hu et al. (2024), and TinyChart Zhang et al. (2024) tackle the reasoning efficiency for high-resolution charts by merging tokens.
@@ 3.1 Architecture
The ChartMoE is based on InternlmXC-v2 Dong et al. (2024) due to the concise LLaVA-like architecture Liu et al. (2023d) and performance on par with GPT-4 on text-image comprehension. The base model includes a vision encoder and a LLM connected by a two-layer MLP. In our ChartMoE  we replace the MLP with an MoE architecture as the connector to leverage diverse prior knowledge.
@@
Large Language Model. Following the baseline, we employ the InternLM2-7B-ChatSFT variant as the LLM $\mathcal{M}^{L}$, implemented as a transformer decoder with a causal attention mask. We concate the visual tokens $\hat{V}:=\{\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{N}\}$ given by MoE connector with the $M$ input text $\mathcal{T}$ tokens $T:=\{t_{1},t_{2},\ldots,t_{M}\}$ to form the input token sequence for the LLM $\mathcal{M}^{L}$. Formally, give the chart $\mathcal{I}$ and instruction $\mathcal{T}$, the output $\mathcal{O}$ of proposed ChartMoE can be formulated as:
@@
Initialization of Expert. Previous approaches initialize expert parameters via 1) Random initialization, which may lead to convergence difficulties during supervised fine-tuning, and 2) Co-upcycling initialization Komatsuzaki et al. (2023), i.e., copy baseline connector parameters to each expert, which may lead to homogenization of experts. ChartMoE introduces a more principled initialization strategy, enabling experts to effectively specialize different components of charts and enhance overall performance over the baseline.
3.2 Alignment Pretraining.
Unlike previous work, we obtain the initialization parameters for the various experts during the different stages of alignment pre-training. Specifically, as illustrated in Fig. 2b, we align expert connectors using three distinct alignment tasks, where only the connector parameters will be updated.
Alignment with Table. ChartAst Meng et al. (2024) first proposes performing chart alignment before supervised fine-tuning, effectively enhancing the model’s ability in numerical extraction and chart understanding. We align charts and tables (in CSV format, shown in Fig. 2a) to improve the model’s capability to extract tabular data.
Alignment with JSON. ChartReformer Yan et al. (2024) proposes using JSON instead of table data for alignment pre-training. JSON contains not only the values of the table but also drawing properties of the chart, such as font format, line type, color style, etc. This task is more difficult than chart-to-table, contributing to attribute-based chart editing and highlighting.
Alignment with Code. Chart editing is primarily achieved by reverse engineering the plotting code and re-rendering the chart Han et al. (2023); Xia et al. (2024). Therefore, we propose the chart-to-code task. Specifically, the code contains precise data and plotting attributes, such as numerical values represented by Python lists and colors represented by hexadecimal codes, as shown in Fig. 2a. This task is more information-rich compared to JSON and Table alignment.
Fig. 3 illustrates how we generate the aligned data in ChartMoE-Align. 1) We filter charts with meta CSV from existing datasets Masry et al. (2022); Methani et al. (2020) and data generated by LLMs Chen et al. (2024). 2) We use a fine-tuned Deplot Liu et al. (2023a) to inverse the plotting attributes following the templates provided by ChartReformer and randomly sample missing attributes from the predefined set. 3) We create code templates for different types of charts and generate plotting code based on the meta CSV and extracted JSON attributes. Note that all values and attributes in the code are explicitly represented. 4) We retain the (table, JSON, code, chart) quadruples that pass Python compilation. Tab. 1 shows the data sources and size in detail.
3.3 Supervised Finetuning.
[…]
@@
High-Quality Knowledge Learning We adopt MMC Liu et al. (2023c) to enhance the ChartMoE’s knowledge. MMC includes a variety of chart types and tasks such as chart-related question answering, translation, extraction, reasoning, and analysis. Considering data quality, we only utilize the MMC-Instruction subset, which has been manually verified. Notice that the quality of instruction data is more important than quantity in this stage.
Chart Specific Annealing Tuning Following Llama-v3.1 Team et al. (2024b), we perform annealing tuning before evaluating mainstream benchmarks. We increase the learning rate and conduct instruction tuning using the training sets of ChartQA Masry et al. (2022) and ChartGemma Masry et al. (2024b) to adjust the query styles and answer formats of these benchmarks.
PoT Chen et al. (2023b) requires the MLLM to generate the variables and operation code rather than producing direct answers. This inference pipeline addresses the mathematical capabilities by employing Python to handle the logical computations, which is the shortcoming of all open-sourced models. With better numerical extraction abilities, PoT can significantly enhance our ChartMoE’s question-answering performance.

### v3

# ChartMoE: Mixture of Diversely Aligned Expert Connector for Chart Understanding

@@ Abstract
Automatic chart understanding is crucial for content comprehension and document parsing. Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in chart understanding through domain-specific alignment and fine-tuning. However, current MLLMs still struggle to provide faithful data and reliable analysis only based on charts. To address it, we propose ChartMoE, which employs the Mixture of Expert (MoE) architecture to replace the traditional linear projector to bridge the modality gap. Specifically, we train several linear connectors through distinct alignment tasks, which are utilized as the foundational initialization parameters for different experts. Additionally, we introduce ChartMoE-Align, a dataset with nearly 1 million chart-table-JSON-code quadruples to conduct three alignment tasks (chart-table/JSON/code). Combined with the vanilla connector, we initialize different experts diversely and adopt high-quality knowledge learning to further refine the MoE connector and LLM parameters. Extensive experiments demonstrate the effectiveness of the MoE connector and our initialization strategy, e.g., ChartMoE improves the accuracy of the previous state-of-the-art from 80.48% to 84.64% on the ChartQA benchmark.
@@ 1 Introduction
Charts serve as a fundamental tool for data visualization, with automated chart interpretation gaining prominence in domains such as text analysis Hoque et al. (2017), scientific research Hsu et al. (2021), and policy-making Wu et al. (2024). Chart understanding is a complex task that demands the identification of visual cues, the comprehension of intricate interactions, and the precise inference of values informed by prior knowledge Huang et al. (2024). Previous work Liu et al. (2023b; a) typically pre-trained on domain-specific charts, which are constrained by limited resources and narrow task focus. In contrast, Multi-modal Large Language Models (MLLMs) Li et al. (2023); Liu et al. (2023d); Bai et al. (2023a); Ye et al. (2023b); Chen et al. (2023); OpenAI (2023) exhibit substantial potential in image comprehension and instruction following. The community has achieved advanced progress by creating chart understanding datasets Liu et al. (2023c); Han et al. (2023); Masry et al. (2024b); Xu et al. (2023) and applying supervised fine-tuning based on well-performed MLLMs Meng et al. (2024); Yan et al. (2024). With the exponential growth of chart data, automated chart interpretation via MLLMs is emerging as a promising avenue.
Recent studies advocate for chart alignment as a foundational step for LLaVA-like MLLMs Liu et al. (2023d); Zhang et al. (2023); Xue et al. (2024), which bridge the visual encoder and LLM through MLP connector. They usually utilize chart-to-table alignment task to train the connector effectively Meng et al. (2024); Yan et al. (2024); Hu et al. (2024). However, tables only provide basic information, such as numerical values and titles, which fail to capture the full range of chart elements. Despite some efforts to align with more informative text Yan et al. (2024), the heavy alignment tasks may lead to the erosion of the connector’s general capabilities, e.g., instruction following and visual counting, which are derived from the pre-training on large-scale visual-language data. To mitigate knowledge forgetting, one intuitive approach is to further tune with its original data, which results in redundant training and computational burden.
@@
Figure 1: Overview and capabilities of ChartMoE: We introduce a MoE architecture connector and provide visualizations of the top-1 expert selection (refer to Fig. 6 and Appendix B for details). ChartMoE can extract highly precise values and provide flexible chart editing through code-based interactions.
In this paper, we try to address these challenges via Mixture of Experts (MoE) architecture Zoph et al. (2022). MoE enhances model capacity by activating a subset of experts through a router. Since the alignment tasks work on the connector, we replace only the MLP projector with MoE while keeping the vision encoder and LLM frozen. Our insight lies in the expert initialization manner. Random initialization can lead to training instability and convergence at sub-optimal points (Fig. 4). Recent co-upcycling initialization Komatsuzaki et al. (2023) addresses this issue by duplicating the vanilla connector parameters across all experts. However, it fails to avoid the dilemma of expert homogenization, where the experts end up with similar functionalities.
In contrast, we attempt to inject distinct prior knowledge into each expert first to tackle these challenges. Unlike natural images, charts can be represented in various text formats, e.g., tables, attribute JSON, and rendering code. As shown in Fig. 12, in addition to chart-table, we introduce chart-JSON alignment to capture detailed elements like color or topological relationships and chart-code alignment to incorporate rendering details such as numerical values, color hex codes, and visual elements interactions (refer to Appendix C). We independently conduct various alignment tasks to capture more diverse chart features and thus obtain three distinct initialization approaches. We also retain the vanilla connector to preserve the capabilities of the MLLM on general tasks effectively.
Building upon the proposed four initialization manners, we introduce ChartMoE, an SFT-based MLLM with MoE connector for chart comprehension and reasoning. Our purpose in applying MoE connector is not to increase model capacity, but rather to improve chart comprehension and diverse representation through alignment tasks, while maintaining performance on other general tasks. Hence, we preserve the original connector parameters as one expert initialization manner. The MoE connector is extremely lightweight, so it adds negligible computational cost during both training and inference. Interestingly, we observe that experts in ChartMoE exhibit distinct visual token preferences, e.g., the vanilla expert favors background tokens while other experts focus more on tokens with legends or numbers (Fig. 6 and Appendix B). Considering that the distribution of visual tokens is naturally imbalanced in chart scenarios, we remove the expert-balanced loss in MoE and obtain further performance gain. Due to the scarcity of rich structural text for chart alignment, we design a pipeline (Fig. 3) to generate nearly 1 million quadruplets chart-table-JSON-code to build the ChartMoE-Align dataset for alignment. We train ChartMoE in 3 stages. First, we initialize experts via the proposed four manners. Then, we conduct high-quality knowledge learning using the MMC instruction Liu et al. (2023c) to train the routing network, expert connectors, and LoRA Hu et al. (2022) modules. Finally, we employ annealing training on ChartQA Masry et al. (2022) and ChartGemma Masry et al. (2024b). ChartMoE achieves state-of-the-art (SOTA) performance and provides more precise numbers and comprehensive attributes of charts (Appendix E). Refer to Appendix D.1 for detailed comparisons with other MoE works w.r.t. motivation, initialization, and complexity. In summary, our contributions are:
@@
    We present ChartMoE for faithful and reasonable chart understanding, with the connector based on Mixture of Expert architecture, to bridge the chart and LLM branches. All experts are initialized based on various alignment training tasks to avoid expert homogenization.
@@
    We introduce ChartMoE-Align, a large-scale dataset with nearly 1 million meticulous chart-table-JSON-code quadruplets for chart alignment pre-training.
@@
    We propose to train ChartMoE with a three-stage training paradigm, including connector alignment pre-training, high-quality knowledge learning, and annealing chart tuning.
@@ 2 Related Work
Multimodal large language models leverages a connector to bridge the gap between large language models Touvron et al. (2023); Radford et al. (2018); Brown et al. (2020); Zhang et al. (2022); Zheng et al. (2023) and vision encoders Radford et al. (2021); Oquab et al. (2023) to enable enriched capabilities of comprehension and instruction following. Approaches such as BLIP2 Li et al. (2023), Flamingo Alayrac et al. (2022), mPLUG-Owl Ye et al. (2023b), and Qwen-VL Bai et al. (2023b) utilize QFormers or Resamplers to align modalities on extensive datasets of image-text pairs. LLaVA Liu et al. (2023d; 2024b) is the pioneering work to extend the instruction tuning paradigm to visual tasks with text-only GPT-4 OpenAI (2023), achieving tremendous performance using a simple MLP without compromising visual information to refine the multimodal alignment. Some works Lin et al. (2023); Tong et al. (2024b; a) explore the combination of various vision encoders, complementarily enhancing visual representations to bolster the fine-grained visual perception of MLLMs. Despite efforts in structural design, training strategies and data quality remain crucial in the advancement of MLLMs.
Chart Reasoning refers to chart analysis, summarization, and etc. Existing methods can be categorized as 1) Two-stage methods use specialized extraction modules to generate intermediary representations of chart information, like tables, which are provided as textual prompts for LLMs. Pix2Struct Lee et al. (2023) aligns markdown data with charts. MatCha Liu et al. (2023b) aligns various data formats (e.g., tables and code) with charts on several downstream tasks. DePlot Liu et al. (2023a) fine-tunes Pix2Struct for table extraction and uses LLMs to process queries based on the extracted data. ChartVLM Xia et al. (2024) employs a discriminator to ascertain the necessity of intervention by LLMs for a given query. 2) End-to-end methods strive to tackle chart reasoning challenges with a unified model. ChartLlama Han et al. (2023) incorporates diverse charts and downstream tasks based on LLaVA Liu et al. (2023d). ChartPaLI Carbune et al. (2024), ChartAst Meng et al. (2024), and MMC Liu et al. (2023c) conduct alignment on table-chart pairs. UReader Ye et al. (2023a) aligns all data with markdown, while mPLUG-Owl2 Ye et al. (2023c) achieves superior performance with high-resolution inputs. ChartThinker Liu et al. (2024c) and DOMINO Wang et al. (2023) propose the CoT Wei et al. (2022) for chart reasoning. LaMenDa Zhuowan et al. (2024) trains MLLMs via step-by-step reasoning QA. ChartReformer Yan et al. (2024) introduces chart-JSON alignment, while OneChart Chen et al. (2024) aligns charts with Python dictionaries. MiniGPT-v2 Chen et al. (2023), Doc-Owl Hu et al. (2024), and TinyChart Zhang et al. (2024) tackle the reasoning efficiency for high-resolution charts by merging tokens.
@@ 3.1 Architecture
The ChartMoE is based on InternlmXC-v2 Dong et al. (2024) due to the concise LLaVA-like architecture Liu et al. (2023d) and performance on par with GPT-4 on text-image comprehension. The base model includes a vision encoder and a LLM connected by a two-layer MLP. ChartMoE replaces the MLP with a MoE architecture as the connector to leverage diverse prior knowledge.
@@
Large Language Model. Following the baseline, we employ the InternLM2-7B-ChatSFT variant as the LLM $\mathcal{M}^{L}$, implemented as a transformer decoder with a causal attention mask. We concate the visual tokens $\hat{V}:=\{\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{N}\}$ given by MoE connector with the $M$ input text $\mathcal{T}$ tokens $T:=\{t_{1},t_{2},\ldots,t_{M}\}$ to form the input token sequence for the LLM $\mathcal{M}^{L}$. Formally, given the chart $\mathcal{I}$ and instruction $\mathcal{T}$, the output $\mathcal{O}$ of proposed ChartMoE can be formulated as:
@@
Initialization of Expert. Previous approaches initialize expert parameters via 1) Random initialization, which may lead to convergence difficulties during supervised fine-tuning, and 2) Co-upcycling initialization Komatsuzaki et al. (2023), i.e., copy baseline connector parameters to each expert, which may lead to homogenization of experts. ChartMoE proposes initializing experts’ parameters through distinct alignment tasks. We eliminate the load-balancing loss typically used in standard MoE architectures to equalize expert activation frequencies, as our initialization approach allows each expert to specialize in its preferred visual tokens, which inherently exhibit biased distributions.
3.2 Alignment Pre-training.
The key insight of ChartMoE is the experts’ initialization parameters from the different alignment pre-training (Fig. 2a). Specifically, as illustrated in Fig. 2b, we align expert connectors using three distinct alignment tasks, where only the connector parameters will be updated. We visualize the visual token preferences of each expert for both chart (Fig. 612) and non-chart (Fig. 11) images.
Alignment with Table. Charts convey key information that can be more precisely expressed in tabular form, and LLMs are particularly adept at processing such structured data. Hence, we introduce a chart-table alignment task, aiming to translate chart content into tabular format. The connector is trained to convert chart information into corresponding CSV tables, thereby improving model performance in numerical data extraction and chart interpretation.
Alignment with JSON. Although tables capture the numerical information from charts, they miss semantic details such as colors, shapes, and fonts. To fill this gap, we propose a chart-JSON alignment task, which represents chart attributes in JSON format. This task requires the connector to focus not only on the numerical data but also on visual and semantic properties. Accurately extracting chart attributes is essential for tasks like chart redrawing and editing.
@@
Alignment with Code. To fully align with charts, we further introduce a chart-code alignment task. Since the underlying drawing code fully defines a chart, this approach enables the connector to convert the chart’s visual tokens into representations in the LLM domain. Notably, we provide the drawing code explicitly, including precise numerical values and rendering attributes, e.g., numbers represented in Python lists and colors in hexadecimal code. Refer to Fig. 13 for more detailed cases. The code enables the model to perform in-depth summarization, analysis, and editing of charts. This expert is significantly more sensitive to the trends and key elements in the charts.
[…]
ChartMoE-Align  generation pipeline. As Fig. 3 illustrates, 1) We filter charts with meta CSV from existing datasets Masry et al. (2022); Methani et al. (2020) and data generated by LLMs Chen et al. (2024). 2) We use a fine-tuned Deplot Liu et al. (2023a) to inverse the plotting attributes following the templates provided by ChartReformer and randomly sample missing attributes from the predefined set. 3) We create code templates for different types of charts and generate plotting code based on the meta CSV and extracted JSON attributes. Note that all values and attributes in the code are explicitly represented. 4) We retain the (table, JSON, code, chart) quadruples that pass compilation. Tab. 1 shows the data sources & size and refer to Appendix C for details.
3.3 Supervised Fine-tuning.
High-Quality Knowledge Learning. We adopt MMC Liu et al. (2023c) to enhance the ChartMoE’s knowledge. MMC includes a variety of chart types and tasks such as chart-related question answering, translation, extraction, reasoning, and analysis. Considering data quality, we only utilize the MMC-Instruction subset, which has been manually verified. Notice that the quality of instruction data is more important than quantity in this stage.
Chart Specific Annealing Tuning. Following Llama-v3.1 Team et al. (2024b), we perform annealing tuning before evaluating mainstream benchmarks. We increase the learning rate and conduct instruction tuning using the training sets of ChartQA and ChartGemma to adjust the query styles and answer formats of these benchmarks.
Program of Thought (PoT) Inference. We require the model to generate the variables and operation code rather than producing direct answers. This inference pipeline addresses the mathematical capabilities by employing Python to handle the logical computations, which is the shortcoming of all open-sourced models. With better numerical extraction abilities, PoT can significantly enhance our ChartMoE’s question-answering performance.